A coffee-cup calorimeter contains 100.0mL of 1.00M HCl at 22.4 degrees celcius. When .243g of Mg metal is added to the acid, the ensuing reaction:

Mg(s) + 2 HCl (aq) --> MgCl2(aq) + H2 (g)

delta Hrxn=?

causes the temperature of the solution to increase to 33.4 degrees celcius. What is the value of delta Hrxn of the reaction? Assume the density of the solution is 1.01 g/ml and that its specific heat is 4.18 J/g x celcius.
Answer #1

heat of reaction = mCpdT

m = 100 ml = 100 ml x 1.01 g/ml = 101 g

Cp = 4.18 J/g.oC

dT = 33.4 - 22.4 = 11.0 oC

we get,

dHrxn = 101 x 4.18 x 11 = 4.64 kJ

           = 4.64 x 24.3/0.243

           = 464 kJ/mol

Thus, Hrxn = 4.64 kJ or 464 kJ/mol